OpenStudy (anonymous):

5/6 + x = 1/-2 - 2/3 x = -1/3 check answer how

OpenStudy (anonymous):

no

OpenStudy (anonymous):

answer is wrong

OpenStudy (anonymous):

x=-1/2-2/3-5/6 x=-3-4-5/6 x=-12/6 x=-2